# Improved finite-size key rates for discrete-modulated continuous variable quantum key distribution under coherent attacks

### v1

@@ Abstract
Continuous variable quantum key distribution (CVQKD) with discrete modulation combines advantages of CVQKD, such as the implementability using readily available technologies, with advantages of discrete variable quantum key distribution, such as easier error correction procedures. We consider a prepare-and-measure CVQKD protocol, where Alice chooses from a set of four coherent states and Bob performs a heterodyne measurement, the result of which is discretised in both key and test rounds. We provide a security proof against coherent attacks in the finite-size regime, and compute the achievable key rate. To this end, we employ the generalised entropy accumulation theorem, as well as recent advances in conic optimisation, yielding improved key rates compared to previous works. At metropolitan distances, our method can provide positive key rates for the order of $10^{8}$ rounds.
@@ I Introduction
Quantum key distribution (QKD) allows two honest parties, conventionally named Alice and Bob, to derive a secret key the security of which is guaranteed by the principles of quantum physics. The first QKD protocols were discrete variable (DV) protocols [1, 2, 3] which typically required the legitimate parties to employ qubits. More recently QKD protocols based on continuous variable (CV) systems, such as squeezed or coherent states [4, 5, 6, 7] have been developed, which can be implemented using readily available optical components and currently existing telecom infrastructure. Hence, continuous variable QKD (CVQKD) protocols present several advantages over their discrete variable QKD (DVQKD) counterparts, including a simpler, affordable implementation that provides optimal scalability, especially at metropolitan scales [8, 9].
@@
Recent results also provide composable, finite-size security proofs against coherent attacks for binary modulated protocols using homodyne and heterodyne detection [34, 35, 36], as well as protocols using four coherent states and heterodyne detection [37]. The former works are based on Azuma’s inequality [38], whereas the latter uses the original version of the entropy accumulation theorem (EAT) [39, 40], applied to an entanglement-based version of the protocol. When applying the proofs to a prepare-and-measure scheme, however, some tricks are required, such as a virtual tomographic measurement on Alice’s side, which lead to decreased non-asymptotic rates.
In this work, we derive a composable security proof against general attacks for a DM CVQKD protocol using four coherent states (also known as 4-PSK protocol) and discretised heterodyne measurements, and provide a finite-size key analysis. The proof is related to the one provided in [37], although it has two main important differences: first, the finite-size analysis is performed via the generalised entropy accumulation theorem (GEAT) [41, 42], which can be directly applied to prepare-and-measure protocols without relying on a virtual tomography, providing improved finite-size secret key rates. Second, for the calculation of the asymptotic key rates, for which we impose a photon number cutoff [28, 37], we use new numerical techniques from conic optimization [43, 44] that provide an enhanced, more reliable performance in contrast with previous approaches based on the Frank-Wolfe algorithm [45]. Our analysis allows us to distill secret key rates for $n\sim 10^{8}$ rounds at metropolitan distances.
@@ II.2 Security definition
A security proof for a QKD protocol then involves two parts: Firstly, it has to be shown that the protocol results in a state that is secret and correct, i.e. close to a fully secret ccq-state. Secrecy and correctness combined are also called soundness. Formally, for $\epsilon^{\mathrm{sou}}>0$, a QKD protocol is said to be $\epsilon^{\mathrm{sou}}$-sound, if it results in a state $\rho^{\mathrm{QKD}}_{ABE}$ after the classical post-processing is performed, such that if we condition on the event ${\Omega_{\text{NA}}}$ of not aborting the protocol it holds
@@ III The protocol
    Alice’s State Preparation: Alice chooses a uniformly random value $x_{j}\in\{0,\ldots,3\}$, corresponding to a random variable $X_{j}$, defining her prepared quantum state $\left|{\varphi_{x_{j}}}\right\rangle_{A^{\prime}}$. She then sends the state $\left|{\varphi_{x_{j}}}\right\rangle_{A^{\prime}}$ to Bob via an insecure quantum channel controlled by Eve. Alice also randomly chooses the nature of the round via a random variable $I_{j}$, taking $I_{j}=0$ for key rounds and $I_{j}=1$ for test rounds (i.e. used for or parameter estimation), whose respective probabilities are $\operatorname{Pr}[I_{j}=0]=p^{\mathrm{K}}$ and $\operatorname{Pr}[I_{j}=1]=p^{\mathrm{PE}}$. For the following discussion, we define separate random variables corresponding to the two types of rounds:
@@
Summarising steps 13, each round $i$ of the protocol creates discrete classical random variables $\hat{X}_{i}$ and $\hat{Z}_{i}$ for key generation, as well as $\tilde{X}_{i}$ and $\tilde{Z}_{i}$ for parameter estimation. We denote $C_{i}=\tilde{X}_{i}\tilde{Z}_{i}$, as the register holding all the classical information used for parameter estimation in round $i$, and $C_{1}^{n}=C_{1}\ldots C_{n}$ for all said registers collected during the execution of the protocol. Let us define Eve’s side information for rounds $1$ to $i$ as $E^{\prime}_{i}$ (so that $E^{\prime}_{i}$ includes all information from previous rounds, $E^{\prime}_{1},\ldots,E^{\prime}_{i-1}$), and we note that Eve has access to registers $I_{i}$ and $\tilde{X}_{i}$ sent by Alice to Bob. All in all, the joint quantum state of Alice, Bob and Eve after $n$ rounds of steps 13 is thus $\sigma^{\mathrm{QKD}}_{\hat{X}_{1}^{n}\hat{Z}_{1}^{n}C_{1}^{n}I_{1}^{n}E^{\prime}_{n}}$. We note also that the round $i+1$ only starts after the $i$-th signal has arrived at Bob’s laboratory. This assumption is needed for ensuring that Eve only possesses one transmitted state at the same time, which is needed for applying the GEAT to prepare-and-measure protocols using the method method of [42] (cf. [42, Condition 3.1]).
Next, Bob performs parameter estimation, which he does by comparing the test results $C_{1}^{n}$ with the probability distribution representing the test results in an ideal, honest implementation. If $C_{1}^{n}$ deviates excessively from the results expected in an ideal run, the protocol is aborted. Let us formalise this notion by considering the string of test results $c_{1}^{n}=(c_{1},...,c_{n})$, where $c_{i}$ takes values from an alphabet $\mathcal{C}$—if round $i$ is a key round, then $c_{i}=(\perp,\perp)$ whereas $c_{i}=(x_{i},z_{i})$ for parameter estimation rounds, for all $x_{i}\in\{0,1,2,3\}$ and $z_{i}\in\{0,1,2,3,4,5\}$. Therefore, we define $\mathcal{C}$ as
@@
for some scalar coefficients $h_{x,z}\in\mathbb{R}$, whose explicit form will be provided later in IV.4. Then, Bob evaluates the affine function $f^{\mathrm{PE}}$ on both probability distributions $\operatorname{freq}_{c_{1}^{n}}$ and $p_{0}$, and he aborts the protocol if and only if $f^{\mathrm{PE}}(\operatorname{freq}_{c_{1}^{n}})<f^{\mathrm{PE}}(p_{0})-\delta^{\mathrm{tol}}$ for a given deviation tolerance $\delta^{\mathrm{tol}}>0$. In other words, the event of passing the parameter estimation without aborting the protocol is defined as
@@ IV Security proof
In this section we show security against coherent attacks. The proof consists of two steps—firstly, in Subsection IV.1 we show the soundness of the protocol and provide a lower bound on the key rate. The soundness proof is based on the GEAT [41, 42] and depends on the choice of a min-tradeoff function of a particular form. Secondly, in Subsection IV.2 we show the completeness of the protocol, i.e. that there is a nonzero probability of it not being aborted. Finally, in Subsection IV.3, we show how a suitable min-tradeoff function can be derived from the numerical approach presented by [45, 29], and build explicitly said function in IV.4.
@@ IV.1 Soundness
$\tilde{\rho}_{ABX}\leq\left(2^{-H^{\epsilon}_{\min}(AX|B)_{\rho}-\delta_{1}}+\delta_{2}\right)\mathbb{1}_{AX}\otimes\sigma_{B}$ (40)
@@
We note that as our channels $\mathcal{M}^{i}_{E_{i-1}\to O_{i}C_{i}E_{i}}$ contain no $R_{i}$ input or outputs, the non-signalling condition given by eq. (1.2) of [41] is trivially satisfied. We can therefore reformulate the GEAT [41, Theorem 4.3] in the following way:
@@ IV.2 Completeness
$1-\operatorname{Pr}[\Omega_{\mathrm{NA}}]_{\sigma^{\mathrm{QKD}}}\big{|}_{\mathrm{hon}}\leq 1-\operatorname{Pr}[\Omega_{\mathrm{PE}}]_{\sigma^{\mathrm{QKD}}}\big{|}_{\mathrm{hon}}+1-\operatorname{Pr}[\Omega_{\mathrm{EC}}]_{\sigma^{\mathrm{QKD}}}\big{|}_{\mathrm{hon}}$ (59)
Let us first consider parameter estimation. Let $p_{0}(x,z)$, a distribution over $(x,z)\in\tilde{\mathcal{C}}$ obtained in an honest implementation, which ios of the form (20) and let $\operatorname{freq}_{c^{n}_{1}}$ be an observed frequency distribution. By the definition of the event $\Omega_{\mathrm{PE}}$, it holds
@@
$\delta^{\mathrm{tol}}\geq 2\sqrt{\log\left(\frac{n}{\epsilon_{\mathrm{PE}}^{c}}\right)\sum_{i=1}^{|\tilde{\mathcal{C}}|}\frac{\gamma_{i}c^{2}_{i}}{n}}+\frac{3D}{n}\log\frac{n}{\epsilon_{\mathrm{PE}}^{c}}$ (65)
@@ IV.3 The Min-Tradeoff Function
We will now derive an affine min-tradeoff function of the form $f(p)=f^{\mathrm{PE}}(p)+\mathrm{const}$, with $f^{\mathrm{PE}}$ defined by eq. (23) [^2] In that case, it holds for all $c_{1}^{n}\in\Omega_{\mathrm{PE}}$ that $f(\operatorname{freq}_{c_{1}^{n}})\geq f(p_{0})-\delta^{\mathrm{tol}}$.. A min-tradeoff function is defined as a function as a function $f:\mathcal{P}_{\mathcal{C}}\to\mathbb{R}$ such that for all $i=1,\ldots,n$ it holds
$\Sigma_{i}(p)=\left\{\left|{\rho}\right\rangle_{E_{i-1}\tilde{E}_{i-1}}\in\mathcal{H}(E_{i-1}\tilde{E}_{i-1}):\left\langle{c}\right|\mathcal{M}^{i}\left(\rho_{E_{i-1}\tilde{E}_{i-1}}\right)\left|{c}\right\rangle\equiv p(c)\right\}.$ (68)
Here $\equiv$ stands for equality for all $c\in\mathcal{C}$. The state $\rho$ can be chosen pure by strong subadditivity as remarked in [39]. As in our protocol the number of key rounds is much larger than the number of test rounds, we make use of the infrequent sampling version of the GEAT. To do so, we decompose the channels $\mathcal{M}^{i}$ as follows
@@
$\tilde{\Sigma}_{i}(\tilde{p})=\left\{\left|{\rho_{E_{i-1}\tilde{E}_{i-1}}}\right\rangle\in\mathcal{H}(E_{i-1}\tilde{E}_{i-1}):\left\langle{c}\right|\mathcal{M}^{i}_{\text{test}}\left(\rho_{E_{i-1}\tilde{E}_{i-1}}\right)\left|{c}\right\rangle\equiv\tilde{p}(c)\right\}.$ (71)
According to Lemma V.5 of [40], we can transform a crossover min-tradeoff function $g$ into a min-tradeoff function $f$ as follows
@@
We can now bound the higher order terms in eqs. (54,55) in terms of $\max(g)$ and $\min(g)$,
@@
$\operatorname{Tr}\left[\left|{x}\right\rangle\left\langle{x}\right|_{A_{i}}\otimes\tilde{R}^{z}_{B_{i}}\left(\tilde{\mathcal{N}}^{i}_{A^{\prime}_{i}E^{\prime}_{i-1}\to B_{i}E^{\prime}_{i}}\left({\psi}_{A_{i}A^{\prime}_{i}}\otimes\rho_{E_{i-1}\tilde{E}_{i-1}}\right)\right)\right]=\tilde{p}(x,z)$ (82)
@@
$\mathcal{M}^{i}_{E_{i-1}\to\hat{Z}_{i}C_{i}E_{i}}(\cdot)=\operatorname{id}_{I_{1}^{i-1}C_{1}^{i-1}}\otimes\widetilde{\mathcal{M}}_{A_{i}B_{i}\to\hat{Z}_{i}C_{i}I_{i}}\circ\tilde{\mathcal{N}}^{i}_{A^{\prime}_{i}E^{\prime}_{i-1}\to B_{i}E^{\prime}_{i}}\left({\psi}_{A_{i}A^{\prime}_{i}}\otimes\cdot\right)$ (83)
@@
$\Sigma^{\prime}_{\hat{E}}(\tilde{p})=\left\{\left|{\omega}\right\rangle_{AB\hat{E}}\in\mathcal{H}(AB\hat{E}):\operatorname{Tr}\left[\left(\left|{x}\right\rangle\left\langle{x}\right|_{A}\otimes\tilde{R}^{z}_{B}\right)\omega_{AB}\right]\equiv\tilde{p}(x,z)\land\omega_{A}=\frac{1}{4}\sum_{i,j=0}^{3}\langle\varphi_{j}|\varphi_{i}\rangle\left|{i}\right\rangle\left\langle{j}\right|_{A}\right\}$ (86)
represents the feasible set (i.e., the set of all states compatible with the constraints imposed by the statistical test). Let us note that the constraint on Alice’s marginal $A$ in (86) can be added since the $A$ system never leaves Alice’s lab, i.e. is not affected by Eve’s attack map. Moreover, since the entanglement of the state (79) cannot be certified from Alice and Bob’s statistics alone, the constraint on Alice’s marginal is necessary to obtain nontrivial results [37]. Following [28, 37], we now consider a coherent version of the map $\widetilde{M}$ defined in (84), corresponding to a coherent version of a round of the protocol. In a coherent version of a round of the protocol the protocol, Alice and Bob’s measurements are performed coherently, i.e. by an isometry writing the result into a quantum register, but not yet depahsing it. This results in a final pure state of the protocol, allowing us to apply [55, Theorem 1], which removes the dependence on Eve’s subsystems. The coherent version of a $\widetilde{M}$ can be expressed in terms of a CP map $\mathcal{G}:AB\to AB\hat{Z}$, given by a single Kraus operator
@@
for $j\in\{0,\ldots,3\}$, and the identity is extended to all registers other than $\hat{Z}$. Using [55, Theorem 1], we can then obtain the following
@@
$\Sigma^{\prime}(\tilde{p})=\left\{\omega_{AB}\in\mathcal{D}(\mathcal{H}_{AB}):\operatorname{Tr}\left[\left(\left|{x}\right\rangle\left\langle{x}\right|_{A}\otimes\tilde{R}_{B}^{z}\right)\omega_{AB}\right]\equiv\tilde{p}(x,z)\land\omega_{A}=\frac{1}{4}\sum_{i,j=0}^{3}\langle\varphi_{j}|\varphi_{i}\rangle\left|{i}\right\rangle\left\langle{j}\right|_{A}\right\},$ (90)

### v2

@@ Abstract
Continuous variable quantum key distribution (CVQKD) with discrete modulation combines advantages of CVQKD, such as the implementability using readily available technologies, with advantages of discrete variable quantum key distribution, such as easier error correction procedures. We consider a prepare-and-measure CVQKD protocol, where Alice chooses from a set of four coherent states and Bob performs a heterodyne measurement, the result of which is discretised in both key and test rounds. We provide a security proof against coherent attacks in the finite-size regime based on the generalised entropy accumulation theorem. We compute the corresponding asymptotic key rate imposing a photon-number cutoff and recent advances in conic optimisation. Our method yields significant improvements on key rates: at metropolitan distances, it provides positive key rates for block sizes of the order of $10^{8}$ rounds, orders of magnitude smaller compared to previous works.
@@ I Introduction
Quantum key distribution (QKD) allows two honest parties, conventionally named Alice and Bob, to derive a secret key the security of which is guaranteed by the principles of quantum physics. The first QKD protocols were discrete variable (DV) protocols [1, 2, 3] which typically required the legitimate parties to employ qubits. More recently, QKD protocols based on continuous variable (CV) states, such as squeezed or coherent states [4, 5, 6, 7] have been developed, which can be implemented using readily available optical components and currently existing telecom infrastructure. Hence, continuous variable QKD (CVQKD) protocols present several advantages over their discrete variable QKD (DVQKD) counterparts, including a simpler, affordable implementation that provides optimal scalability, especially at metropolitan scales [8, 9].
@@
Recent results also provide composable, finite-size security proofs against coherent attacks for binary modulated protocols using homodyne and heterodyne detection [34, 35, 36], as well as protocols using four coherent states and heterodyne detection [37]. The former works are based on Azuma’s inequality [38], whereas the latter uses the original version of the entropy accumulation theorem (EAT) [39, 40], applied to an entanglement-based version of the protocol. When applying the proofs to a prepare-and-measure scheme, however, some challenges arise, requiring a virtual tomographic measurement on Alice’s side, which lead to decreased non-asymptotic rates.
In this work, we derive a composable security proof against general attacks for a DM CVQKD protocol using four coherent states (also known as 4-PSK protocol) and discretised heterodyne measurements, and provide a finite-size key analysis. The proof is related to the one provided in [37], although it has two main important differences: first, the finite-size analysis is performed via the generalised entropy accumulation theorem (GEAT) [41, 42], which can be directly applied to prepare-and-measure protocols without relying on a virtual tomography, providing improved finite-size secret key rates. Second, for the calculation of the asymptotic key rates, for which we impose a photon number cutoff [28, 37], we use new numerical techniques from conic optimization [43, 44, 45] that provide an enhanced, more reliable performance in contrast with previous approaches based on the Frank-Wolfe algorithm [46]. Our analysis allows us to distill secret key rates for $n\sim 10^{8}$ rounds at metropolitan distances.
@@ II.2 Security definition
A security proof for a QKD protocol then involves two parts: Firstly, it has to be shown that the protocol results in a state that is secret and correct, i.e. close to a fully secret ccq-state. Secrecy and correctness combined are also called soundness. Formally, for $\epsilon^{\mathrm{sou}}>0$, a QKD protocol is said to be $\epsilon^{\mathrm{sou}}$-sound if it results in a state $\rho^{\mathrm{QKD}}_{ABE}$ after the classical post-processing is performed such that, if we condition on the event ${\Omega_{\text{NA}}}$ of not aborting the protocol, it holds
@@ III The protocol
    Alice’s State Preparation: Alice chooses a uniformly random value $x_{j}\in\{0,\ldots,3\}$, corresponding to a random variable $X_{j}$, defining her prepared quantum state $\left|{\varphi_{x_{j}}}\right\rangle_{A^{\prime}}$. She then sends the state $\left|{\varphi_{x_{j}}}\right\rangle_{A^{\prime}}$ to Bob via an insecure quantum channel controlled by Eve. Alice also randomly chooses the kind of round via a random variable $I_{j}$, taking $I_{j}=0$ for key rounds and $I_{j}=1$ for test rounds (i.e. used for or parameter estimation), whose respective probabilities are $\operatorname{Pr}[I_{j}=0]=p^{\mathrm{K}}$ and $\operatorname{Pr}[I_{j}=1]=p^{\mathrm{PE}}$. For the following discussion, we define separate random variables corresponding to the two types of rounds:
@@
Summarising steps 13, each round $i$ of the protocol creates discrete classical random variables $\hat{X}_{i}$ and $\hat{Z}_{i}$ for key generation, as well as $\tilde{X}_{i}$ and $\tilde{Z}_{i}$ for parameter estimation. We denote $C_{i}=\tilde{X}_{i}\tilde{Z}_{i}$, as the register holding all the classical information used for parameter estimation in round $i$, and $C_{1}^{n}=C_{1}\ldots C_{n}$ for all said registers collected during the execution of the protocol. Let us define Eve’s side information for rounds $1$ to $i$ as $E^{\prime}_{i}$ (so that $E^{\prime}_{i}$ includes all information from previous rounds, $E^{\prime}_{1},\ldots,E^{\prime}_{i-1}$), and we note that Eve has access to registers $I_{i}$ and $\tilde{X}_{i}$ sent by Alice to Bob. All in all, the joint quantum state of Alice, Bob and Eve after $n$ rounds of steps 13 is thus $\sigma^{\mathrm{QKD}}_{\hat{X}_{1}^{n}\hat{Z}_{1}^{n}C_{1}^{n}I_{1}^{n}E^{\prime}_{n}}$. We note also that the round $i+1$ only starts after the $i$-th signal has arrived at Bob’s laboratory. This is used to ensure that Eve only possesses one transmitted state at the same time, which is needed for applying the GEAT to prepare-and-measure protocols using the method of [42] (cf. [42, Condition 3.1]).
Next, Bob performs parameter estimation, which he does by comparing the test results $C_{1}^{n}$ with the probability distribution representing the test results in an ideal, honest implementation. If $C_{1}^{n}$ deviates excessively from the results expected in an ideal run, the protocol is aborted. Let us formalise this notion by considering the string of test results $c_{1}^{n}=(c_{1},...,c_{n})$, where $c_{i}$ takes values from an alphabet $\mathcal{C}$—if round $i$ is a key round, then $c_{i}=(\perp,\perp)$ whereas $c_{i}=(x_{i},z_{i})$ for parameter estimation rounds, for all $x_{i}\in\{0,1,2,3\}$ and $z_{i}\in\{0,1,2,3,4,5\}$. Thus, we define $\mathcal{C}$ as
@@
for some scalar coefficients $h_{x,z}\in\mathbb{R}$, whose explicit form will be provided later in IV.4. We will also see that the function $f^{\mathrm{PE}}$ is close to what, in the context of GEAT, is referred to as a min-tradeoff function (see eq. (53) for an exact definition). Bob then evaluates the affine function $f^{\mathrm{PE}}$ on both probability distributions $\operatorname{freq}_{c_{1}^{n}}$ and $p_{0}$, and he aborts the protocol if and only if $f^{\mathrm{PE}}(\operatorname{freq}_{c_{1}^{n}})<f^{\mathrm{PE}}(p_{0})-\delta^{\mathrm{tol}}$ for a given deviation tolerance $\delta^{\mathrm{tol}}>0$. In other words, the event of passing the parameter estimation without aborting the protocol is defined as
@@ IV Security proof
In this section we show security against coherent attacks. The proof consists of two steps—firstly, in Subsection IV.1 we show the soundness of the protocol and provide a lower bound on the key rate. The soundness proof is based on the GEAT [41, 42], a tool allowing us to lower bound the amount of key obtainable in the presence of coherent attacks by reducing the problem to the case of only collective attacks. Secondly, in Subsection IV.2 we show the completeness of the protocol, i.e. that there is a nonzero probability of it not being aborted. Finally, in Subsection IV.3, we use techniques based on [46, 29] to numerically compute the collective attack bounds.
@@ IV.1 Soundness
$\tilde{\rho}_{ABX}\leq\left(2^{-H^{\epsilon}_{\min}(AX|B)_{\rho}-\delta_{1}}+\delta_{2}\right)\mathbb{1}_{AX}\otimes\sigma_{B}.$ (40)
@@
We note that as our channels $\mathcal{M}^{i}_{E_{i-1}\to O_{i}C_{i}E_{i}}$ contain no $R_{i}$ input or outputs, the non-signalling condition given by eq. (1.2) of [41] is trivially satisfied. In order to apply the GEAT, we further need to introduce the concept of a min-tradeoff function. A min-tradeoff function $f:\mathcal{P}_{\mathcal{C}}\to\mathbb{R}$ is defined as a real valued function on the distributions of the alphabet $\mathcal{C}$, which we have defined in eq. (18), such that for all $i=1,\ldots,n$ it holds
$\Sigma_{i}(p)=\left\{\left|{\rho}\right\rangle_{E_{i-1}\tilde{E}_{i-1}}\in\mathcal{H}(E_{i-1}\tilde{E}_{i-1}):\operatorname{Tr}\left[\left|{c}\right\rangle\left\langle{c}\right|_{C_{i}}\otimes\operatorname{id}_{O_{i}E_{i}}\mathcal{M}^{i}\left(\rho_{E_{i-1}\tilde{E}_{i-1}}\right)\right]=p(c)\;\forall c\in\mathcal{C}\right\}.$ (54)
The state $\rho$ can be chosen pure by strong subadditivity as remarked in [39]. We can now reformulate the GEAT [41, Theorem 4.3] in the following way:
@@ IV.2 Completeness
$1-\operatorname{Pr}[\Omega_{\mathrm{NA}}]_{\sigma^{\mathrm{QKD}}}\big{|}_{\mathrm{hon}}\leq 1-\operatorname{Pr}[\Omega_{\mathrm{PE}}]_{\sigma^{\mathrm{QKD}}}\big{|}_{\mathrm{hon}}+1-\operatorname{Pr}[\Omega_{\mathrm{EC}}]_{\sigma^{\mathrm{QKD}}}\big{|}_{\mathrm{hon}}.$ (61)
Let us first consider parameter estimation. Let $p_{0}(x,z)$, a distribution over $(x,z)\in\tilde{\mathcal{C}}$ obtained in an honest implementation, which is of the form (20) and let $\operatorname{freq}_{c^{n}_{1}}$ be an observed frequency distribution. By the definition of the event $\Omega_{\mathrm{PE}}$, it holds
@@
$\delta^{\mathrm{tol}}\geq 2\sqrt{\log\left(\frac{n}{\epsilon_{\mathrm{PE}}^{c}}\right)\sum_{i=1}^{|\tilde{\mathcal{C}}|}\frac{\gamma_{i}c^{2}_{i}}{n}}+\frac{3D}{n}\log\frac{n}{\epsilon_{\mathrm{PE}}^{c}},$ (67)
@@ IV.3 The Min-Tradeoff Function
We will now derive an affine min-tradeoff function of the form $f(p)=f^{\mathrm{PE}}(p)+\mathrm{const}$, with $f^{\mathrm{PE}}$ defined by eq. (23) [^2] In that case, it holds for all $c_{1}^{n}\in\Omega_{\mathrm{PE}}$ that $f(\operatorname{freq}_{c_{1}^{n}})\geq f(p_{0})-\delta^{\mathrm{tol}}$.. As in our protocol the number of key rounds is much larger than the number of test rounds, we make use of the infrequent sampling version of the GEAT. To do so, we decompose the channels $\mathcal{M}^{i}$ as follows
@@
$\tilde{\Sigma}_{i}(\tilde{p})=\left\{\left|{\rho_{E_{i-1}\tilde{E}_{i-1}}}\right\rangle\in\mathcal{H}(E_{i-1}\tilde{E}_{i-1}):\operatorname{Tr}\left[\left|{c}\right\rangle\left\langle{c}\right|_{C_{i}}\operatorname{id}_{O_{i}E_{i}}\mathcal{M}^{i}_{\text{test}}\left(\rho_{E_{i-1}\tilde{E}_{i-1}}\right)\right]=\;\tilde{p}(c)\forall c\in\tilde{\mathcal{C}}\right\}.$ (71)
Note that this constraint set differs from the one given by eq. (54) in that the constraints are only in terms of $\mathcal{M}^{i}_{\text{test}}$. According to Lemma V.5 of [40], we can transform a crossover min-tradeoff function $g$ into a min-tradeoff function $f$ as follows
@@
We can now bound the higher order terms in eqs. (56,57) in terms of $\max(g)$ and $\min(g)$,
@@
$\operatorname{Tr}\left[\left|{x}\right\rangle\left\langle{x}\right|_{A_{i}}\otimes\tilde{R}^{z}_{B_{i}}\left(\tilde{\mathcal{N}}^{i}_{A^{\prime}_{i}E^{\prime}_{i-1}\to B_{i}E^{\prime}_{i}}\left({\psi}_{A_{i}A^{\prime}_{i}}\otimes\rho_{E_{i-1}\tilde{E}_{i-1}}\right)\right)\right]=\tilde{p}(x,z),$ (82)
@@
$\mathcal{M}^{i}_{E_{i-1}\to\hat{Z}_{i}C_{i}E_{i}}(\cdot)=\operatorname{id}_{I_{1}^{i-1}C_{1}^{i-1}}\otimes\widetilde{\mathcal{M}}_{A_{i}B_{i}\to\hat{Z}_{i}C_{i}I_{i}}\circ\tilde{\mathcal{N}}^{i}_{A^{\prime}_{i}E^{\prime}_{i-1}\to B_{i}E^{\prime}_{i}}\left({\psi}_{A_{i}A^{\prime}_{i}}\otimes\cdot\right),$ (83)
@@
$\Sigma^{\prime}_{\hat{E}}(\tilde{p})=\left\{\left|{\omega}\right\rangle_{AB\hat{E}}\in\mathcal{H}(AB\hat{E}):\operatorname{Tr}\left[\left(\left|{x}\right\rangle\left\langle{x}\right|_{A}\otimes\tilde{R}^{z}_{B}\right)\omega_{AB}\right]=\tilde{p}(x,z)\;\forall(x,z)\in\tilde{\mathcal{C}}\land\omega_{A}=\frac{1}{4}\sum_{i,j=0}^{3}\langle\varphi_{j}|\varphi_{i}\rangle\left|{i}\right\rangle\left\langle{j}\right|_{A}\right\}$ (86)
represents the feasible set (i.e., the set of all states compatible with the constraints imposed by the statistical test). Let us note that the constraint on Alice’s marginal $A$ in (86) can be added since the $A$ system never leaves Alice’s lab, i.e. is not affected by Eve’s attack map. Moreover, since the entanglement of the state (79) cannot be certified from Alice and Bob’s statistics alone, the constraint on Alice’s marginal is necessary to obtain nontrivial results [37]. Following [28, 37], we now consider a coherent version of the map $\widetilde{M}$ defined in (84), corresponding to a coherent version of a round of the protocol. In a coherent version of a round of the protocol, Alice and Bob’s measurements are performed coherently, i.e. by an isometry writing the result into a quantum register, but not yet depahsing it. This results in a final pure state of the protocol, allowing us to apply [56, Theorem 1], which removes the dependence on Eve’s subsystems. The coherent version of a $\widetilde{M}$ can be expressed in terms of a CP map $\mathcal{G}:AB\to AB\hat{Z}$, given by a single Kraus operator
@@
for $j\in\{0,\ldots,3\}$, and the identity is extended to all registers other than $\hat{Z}$. Using [56, Theorem 1], we can then obtain the following
@@
$\Sigma^{\prime}(\tilde{p})=\left\{\omega_{AB}\in\mathcal{D}(\mathcal{H}_{AB}):\operatorname{Tr}\left[\left(\left|{x}\right\rangle\left\langle{x}\right|_{A}\otimes\tilde{R}_{B}^{z}\right)\omega_{AB}\right]=\tilde{p}(x,z)\;\forall(x,z)\in\tilde{\mathcal{C}}\land\omega_{A}=\frac{1}{4}\sum_{i,j=0}^{3}\langle\varphi_{j}|\varphi_{i}\rangle\left|{i}\right\rangle\left\langle{j}\right|_{A}\right\},$ (90)